OpenStudy (anonymous):

use an identity to solve each equation on interval [0,2pi) cos2x=sinx+1

OpenStudy (abb0t):

subtract sin(x)+1 from both sides first.

OpenStudy (abb0t):

then use your trig identity to change cos(2x) into one of the \(sin^2(x)\) half trig formulas or whatveer they are called.

OpenStudy (abb0t):

If you need more help, don't hesistate to ask.

OpenStudy (aravindg):

\[1-\cos 2x=2\sin^2 x\]\[\cos 2x=1-2\sin^2x\]Replace \(\cos 2x \) with \(1-2\sin^2 x\) in the question and solve.